Find the center and vertices of the ellipse.
4x^2+y^2=4
a.center:(-2,2) vertices:(-1,0),(1,0)
b.center:(-2,2) vertices:(-1,-2),(1,2)
c.center:(0,0) vertices:(-2,0),(2,0)
d.center:(0,0) vertices:(-2,-1),(2,1)
e.center:(0,0) vertices:(0,-2),(0,2)
Answer by lwsshak3(11628) 
 Find the center and vertices of the ellipse.
4x^2+y^2=4
**
4x^2+y^2=4
divide by 4
x^2+y^2/4=1
This is an ellipse with vertical major axis.
Standard form for ellipse with vertical major axis: (x-h)^2/b^2+(y-k)^2/a^2=1 a>b
For given ellipse:
Center: (0,0)
a^2=4
a=2
length of major axis=2a=4
Vertices: (0,0±a) = (0,0±2) or (0,-2) and (0,2)
ans:
e.center:(0,0) vertices:(0,-2),(0,2)
| |
|